OpenStudy (anonymous):

Write an equation of the line that is parallel to -x + y = 5 and passes through the point (2, -5).

OpenStudy (nottim):

closed? already?

OpenStudy (anonymous):

Okay Im working on the problem now . do you still need the answer?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

Okay , No problem , ill tell you when im done.

OpenStudy (anonymous):

ok thanks :)

OpenStudy (anonymous):

-x + y = 5 y=x+5 here slope m1 = 1.......(1) slope of line that is parallel to the (1) is m2=1 equation=> y+5 =1(x-2) y+5 =x-2 x-y-7=0 There you go !

OpenStudy (anonymous):

Dont forget to medal me (: lol

OpenStudy (anonymous):

ok;)